# Fanxu Meng1,2,3, Zhaohui Wang1, Muhan Zhang1,3

### v3

@@ 5 Experiments
The experiments were conducted on the NVIDIA A800-SXM4(80G) GPU. In our experiments, we adopt the Alpaca [52] implementation strategy, using the AdamW optimizer with a batch size of 128, a learning rate of 2e-5, cosine annealing schedules, and a warmup ratio of 0.03, without any weight decay. As discussed in Section B.3 of QLoRA [10], we compute the loss using only the responses from the instruction-following datasets. We ensure lora_alpha is always equal to lora_r, set lora_dropout to 0, and incorporate the adapters into all linear layers of the base model. We utilize the Float32 computation type for both the base model and the adapter in LoRA and PiSSA. For QLoRA, LoftQ, and QPiSSA, we use 4-bit NormalFloat [10] for the base model and Float32 for the adapter. BFloat16 [53] is used for full parameter fine-tuning to save the resources (see Appendix D).
@@ 5.1 Evaluating the Performance of PiSSA on both NLG and NLU Tasks
Table 1 presents a comparative evaluation of fine-tuning strategies PiSSA, LoRA, and full parameter fine-tuning on natural language generation (NLG) tasks. We fine-tuned LLaMA 2-7B [54], Mistral-7B-v0.1 [55], and Gemma-7B [56] on the MetaMathQA dataset [2] to assess their mathematical problem-solving capabilities on the GSM8K [57] and MATH [2] validation sets. Additionally, the models were fine-tuned on the CodeFeedback dataset [58] and evaluated for coding proficiency using the HumanEval [59] and MBPP [60] datasets. Furthermore, the models were trained on the WizardLM-Evol-Instruct dataset [7] and tested for conversational abilities on the MT-Bench dataset [6]. All experiments were conducted using subsets containing 100K data points and were trained for only one epoch to reduce training overhead.
Table 1: Comparing PiSSA and LoRA on NLG tasks
Model Strategy Trainable GSM8K MATH HumanEval MBPP MT-Bench
Parameters
LLaMA 2-7B Full FT 6738M 49.05 7.22 21.34 35.59 4.91
LoRA 320M 42.30 5.50 18.29 35.34 4.58
PiSSA 320M 53.07 7.44 21.95 37.09 4.87
Mistral-7B Full FT 7242M 67.02 18.6 45.12 51.38 4.95
LoRA 168M 67.70 19.68 43.90 58.39 4.90
PiSSA 168M 72.86 21.54 46.95 62.66 5.34
Gemma-7B Full FT 8538M 71.34 22.74 46.95 55.64 5.40
LoRA 200M 74.90 31.28 53.66 65.41 4.98
PiSSA 200M 77.94 31.94 54.27 66.17 5.64
@@
As shown in Table 1, across all models and tasks, fine-tuning with PiSSA consistently surpasses the performance of fine-tuning with LoRA. For instance, fine-tuning the LLaMA, Mistral, and Gemma with PiSSA results in performance improvements of 10.77%, 5.26%, and 3.04% respectively in mathematical tasks. In coding tasks, the improvements are 20%, 6.95%, and 1.14% respectively. In MT-Bench, we observe enhancements of 6.33%, 8.98%, and 13.25%. Notably, using PiSSA with only 2.3% of Gemma’s trainable parameters outperforms full parameter fine-tuning by 15.59% in coding tasks. Further experiments demonstrated that this improvement is robust across various amounts of training data and epochs (Section 5.2), including both 4-bit and full precision (Section 5.3), different model sizes and types (Section 5.4), and varying proportions of trainable parameters (Section 5.5).
Table 2: Comparing PiSSA and LoRA on NLU tasks. Details can be found in Appendix  I.
Method Parameters MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
RoBERTa-large (355M)
Full FT 355M 90.2 96.4 90.9 68.0 94.7 92.2 86.6 91.5
LoRA 1.84M 90.6 96.2 90.9 68.2 94.9 91.6 87.4 92.6
PiSSA 1.84M 90.7 96.7 91.9 69.0 95.1 91.6 91.0 92.9
DeBERTa-v3-base (184M)
Full FT 184M 89.90 95.63 89.46 69.19 94.03 92.40 83.75 91.60
LoRA 1.33M 90.65 94.95 89.95 69.82 93.87 91.99 85.20 91.60
PiSSA 1.33M 90.43 95.87 91.67 72.64 94.29 92.26 87.00 91.88
@@
We also evaluate PiSSA’s natural language understanding (NLU) capability on the GLUE benchmark [61] with RoBERTa-large [62] and DeBERTa-v3-base [63]. Table 2 presents the results across 8 tasks conducted with the two base models. PiSSA outperforms LoRA on 14 out of 16 experimental settings, and achieves the same performance on QQP + RoBERTa. The single exception arises from the DeBERTa-based model on the MNLI dataset. Upon reviewing the training loss, we observed that PiSSA’s average loss of $0.17$ was lower than LoRA’s $0.24$ in the final epoch. This indicates that the fitting ability of PiSSA remains stronger than that of LoRA.
@@ 5.2 Experiments using Full Data and More Epochs
In this section, we finetune LLaMA 2-7B model on the complete MetaMathQA-395K dataset for 3 epochs to ensure thorough saturation. The training loss and gradient norms is visualized to demonstrate quicker convergence and evaluated on the GSM8K dataset every 1000 steps to demonstrate superior performance of PiSSA compared to LoRA. The results are depicted in Figure 4. Additionally, similar comparisons on Mistral-7B and Gemma-7B are detailed in Appendix G.
@@
Figure 4: The loss, grad norm, and evaluation accuracy over the training steps of LoRA (indicated in blue), PiSSA (in orange), and full parameter fine-tuning (in green).
@@ 5.3 Conducting 4-bit Quantization Experiments
In Table 3, PiSSA reduces the quantization error by about 20% compared to directly quantizing the base model. The reduction is more significant for lower-rank matrices. For instance, in the LLaMA-3-70B [64], all “Key” projection layers see a reduction of 49%. The results in Table 3 validate that QLoRA, discussed in Section 4, does not reduce quantization error. In contrast, PiSSA significantly outperforms LoftQ in reducing quantization error, as further discussed in Appendix F.
Besides reducing the quantization error, we expect QPiSSA to also converge faster than QLoRA and LoftQ. We train LLaMA 3-8B with LoRA/QLoRA, PiSSA/QPiSSA, LoftQ, and full fine-tuning on MetaMathQA-395K for 3 epochs and record the loss, grad norm, and accuracy on GSM8K.
@@ 5.4 Experiments Across Various Sizes and Types of Models
In this section, we compare (Q)PiSSA and (Q)LoRA across 9 models, ranging from 7-70B parameters, including LLaMA 2-7/13B [54], LLaMA-3-8/70B [64], Mistral-7B [55], Gemma-7B [56], and Qwen1.5-7B [65], Yi-1.5-34B [66] and MoE models: DeepSeek-MoE-16B [67] and Mixtral-8x7B [68]. These models were fine-tuned on the MetaMathQA-100K and CodeFeedback-100K dataset and evaluated on the GSM8K and HumanEval. DeepSeek-MoE-16B, Mixtral-8x7B, Yi-1.5-34B, and LLaMA-3-70B were fine-tuned with QPiSSA and QLoRA, while the other models were using PiSSA and LoRA. From Figure 6, it can be observed that (Q)PiSSA, compared to (Q)LoRA, shows improved accuracy across various sizes and types of models, demonstrating its consistent advantage over (Q)LoRA.
@@ 5.5 Experiments on Various Ranks
This section explores the impact of incrementally increasing the rank of PiSSA/QPiSSA and LoRA/QLoRA from 1 to 128, aiming to determine whether PiSSA/QPiSSA consistently outperforms LoRA/QLoRA under different ranks. The training is conducted using the MetaMathQA-100K dataset for 1 epoch, while the validation is performed on the GSM8K and MATH datasets. The outcomes of these experiments are depicted in Figure 7, with additional results presented in Appendix H.
@@ 6 Conclusion
This paper presents a PEFT technique that applies singular value decomposition (SVD) to the weight matrix of pre-trained models. The principal components obtained from the SVD are used to initialize a low-rank adapter named PiSSA, while the residual components are kept frozen, in order to achieve effective fine-tuning and parameter efficiency at the same time. Through extensive experiments, we found that PiSSA and its 4-bit quantization version QPiSSA significantly outperform LoRA and QLoRA in both NLG and NLU tasks, across different training steps, various model sizes and types, and under various amount of trainable parameters. PiSSA provides a novel direction for research in PEFT by identifying and fine-tuning the principal components within the model, analogous to slicing and re-baking the richest slice of a pizza. As PiSSA shares the same architecture as LoRA, it can be seamlessly used in existing LoRA pipelines as an efficient alternative initialization method.
@@ 7 Limitation
There are still some questions with PiSSA not addressed in this paper: 1) Besides language models, can PiSSA also be adapted to convolutional layers and enhance the performance of vision tasks? 2) Can PiSSA also benefit from some improvements to LoRA, such as AdaLoRA [69] and DyLoRA [42] which adaptively adjust the rank? 3) Can we provide more theoretical explanations for the advantages of PiSSA over LoRA? We are actively exploring these questions. Nevertheless, we are excited to see the huge potential of PiSSA already demonstrated in existing experiments and look forward to more tests and suggestions from the community.
Appendix A Conductive Experiments on Various SVD Components
@@ Appendix D Comparison of Fine-Tuning in BF16 and FP32 Precision
Appendix E Reducing Quantization Error through QPiSSA with Multiple Iteration of SVD
@@
Appendix F Comparing the Quantization Error of QLoRA, LoftQ and QPiSSA
@@
Appendix G Evaluating PiSSA on Mixtral and Gemma with More Training Steps
This is the supplement for Section 5.2. We applied PiSSA, LoRA, and full parameter fine-tuning on the full MetaMathQA-395K dataset using Mistral-7B and Gemma-7B models, training for 3 epochs. Figures 11 and 12 display the training loss, gradient norm, and evaluation accuracy on GSM8K.

### v4

@@ 5 Experiments
The experiments were conducted on the NVIDIA A800-SXM4(80G) GPU. In our experiments, we adopt the Alpaca alpaca implementation strategy, using the AdamW optimizer with a batch size of 128, a learning rate of 2e-5, cosine annealing schedules, and a warmup ratio of 0.03, without any weight decay. As discussed in Section B.3 of QLoRA dettmers2024qlora , we compute the loss using only the responses from the instruction-following datasets. We ensure lora_alpha is always equal to lora_r, set lora_dropout to 0, and incorporate the adapters into all linear layers of the base model. We utilize the Float32 computation type for both the base model and the adapter in LoRA and PiSSA. For QLoRA, LoftQ, and QPiSSA, we use 4-bit NormalFloat dettmers2024qlora for the base model and Float32 for the adapter. BFloat16 wang2019bfloat16 is used for full parameter fine-tuning to save the resources (see Appendix D).
@@ 5.1 Evaluating the Performance of PiSSA on both NLG and NLU Tasks
We begin by comparing PiSSA, LoRA, and full-parameter fine-tuning on natural language generation (NLG) tasks. We fine-tuned LLaMA 2-7B touvron2023llama , Mistral-7B-v0.1 jiang2023mistral , and Gemma-7B team2024gemma on the MetaMathQA dataset yu2023metamath to assess their mathematical problem-solving capabilities on the GSM8K cobbe2021gsm8k and MATH hendrycks2021measuring validation sets. Additionally, the models were fine-tuned on the CodeFeedback dataset zheng2024opencodeinterpreter and evaluated for coding proficiency using the HumanEval chen2021evaluating and MBPP austin2021program datasets. Furthermore, the models were trained on the WizardLM-Evol-Instruct dataset xu2023wizardlm and tested for conversational abilities on the MT-Bench dataset zheng2024judging . All experiments were conducted using subsets containing 100K data points and were trained for only one epoch to reduce training overhead.
Table 2: Comparison of PiSSA and LoRA on NLG tasks, with results averaged over three runs and reported with standard deviations.
Model Strategy GSM8K MATH HumanEval MBPP MT-Bench
LLaMA 2-7B Full FT 49.13±0.21 7.29±0.22 21.20±0.30 35.59±0.25 4.91±0.01
LoRA(gaussian) 42.85±0.12 5.50±0.33 18.35±0.31 35.50±0.14 4.59±0.07
LoRA(kaiming) 43.23±0.64 5.90±0.16 18.21±0.45 35.47±0.37 4.56±0.04
PiSSA 53.22±0.55 7.47±0.34 21.92±0.38 37.24±0.63 4.88±0.03
Mistral-7B Full FT 69.91±0.25 18.64±0.35 45.31±0.14 51.46±0.13 4.95±0.05
LoRA(gaussian) 69.50±0.42 20.08±0.20 43.78±0.34 58.46±0.37 4.90±0.05
LoRA(kaiming) 69.40±0.25 19.99±0.44 43.74±0.14 58.39±0.02 4.93±0.05
PiSSA 73.31±0.23 23.12±0.52 46.88±0.25 62.55±0.58 5.34±0.04
Gemma-7B Full FT 72.09±0.32 22.71±0.34 47.02±0.27 55.67±0.50 5.40±0.12
LoRA(gaussian) 75.11±0.64 30.41±0.48 53.70±0.23 65.58±0.29 4.98±0.02
LoRA(kaiming) 74.53±0.47 29.90±0.16 53.57±0.27 65.25±0.29 4.97±0.09
PiSSA 77.78±0.32 31.33±0.33 54.31±0.28 66.17±0.43 5.64±0.10
@@
As shown in Table 2, across all models and tasks, fine-tuning with PiSSA consistently surpasses the performance of fine-tuning with LoRA. Further experiments demonstrated that this improvement is robust across various amounts of training data and epochs (Section 5.2), including both 4-bit and full precision (Section 5.3), different model sizes and types (Section 5.4), and varying proportions of trainable parameters (Section 5.5).
Table 3: Comparison of PiSSA and LoRA on NLU tasks. LoRAG and LoRAK denote LoRA with Gaussian and Kaiming initialization for $B$, respectively. Results for full fine-tuning, BitFit zaken2021bitfit , HAdapter houlsby2019parameter , PAdapter pfeiffer2020adapterfusion , LoRAG hu2021lora and AdaLoRA are from AdaLoRA zhang2023adaptive , averaged over five runs. Remaining methods are averaged over three runs, with details in Appendix L.
Method Params MNLI SST2 MRPC CoLA QNLI QQP RTE STSB ALL
Full FT 184M 89.90 95.63 89.46 69.19 94.03 92.40 83.75 91.60 88.25
BitFit 0.1M 89.37 94.84 87.75 66.96 92.24 88.41 78.70 91.35 86.20
HAdapter 1.22M 90.13 95.53 89.95 68.64 94.11 91.91 84.48 91.48 88.28
PAdapter 1.18M 90.33 95.61 89.46 68.77 94.29 92.04 85.20 91.54 88.41
LoRAG 1.33M 90.65 94.95 89.95 69.82 93.87 91.99 85.20 91.60 88.50
LoRAK 1.33M 89.96 95.64 90.28 70.69 93.84 92.03 84.84 91.68 88.62
DoRA 1.27M 90.29 95.79 90.93 70.85 94.10 92.07 86.04 91.79 88.98
AdaLoRA 1.27M 90.76 96.10 90.69 71.45 94.55 92.23 88.09 91.84 89.46
PiSSA 1.33M 90.37 96.22 91.50 73.12 94.43 92.33 88.69 92.00 89.83
@@
We also evaluate PiSSA’s natural language understanding (NLU) capability on the GLUE benchmark wang2018glue with DeBERTa-v3-base he2021debertav3 . Table 3 presents the results across 8 tasks. PiSSA outperforms LoRA in 7 out of 8 NLU tasks, achieving an overall average improvement of 1.21%. Upon reviewing the training loss on the exceptional dataset, MNLI, we observed that PiSSA’s average loss of $0.17$ was lower than LoRA’s $0.24$ in the final epoch. This indicates that the fitting ability of PiSSA remains stronger than that of LoRA.
@@ 5.2 Experiments using Full Data and More Epochs
In this section, we finetune LLaMA 2-7B model on the complete MetaMathQA-395K dataset for 3 epochs to ensure thorough saturation. The training loss and gradient norms is visualized to demonstrate quicker convergence and evaluated on the GSM8K dataset every 1000 steps to demonstrate superior performance of PiSSA compared to LoRA. The results are depicted in Figure 4. Additionally, similar comparisons on Mistral-7B and Gemma-7B are detailed in Appendix J.
@@
Figure 4: The loss, grad norm, and evaluation accuracy over the training steps of LoRA (indicated in blue), PiSSA (in orange), and full parameter fine-tuning (in red).
@@ 5.3 Conducting 4-bit Quantization Experiments
In Table 4, PiSSA reduces the quantization error by about 20% compared to directly quantizing the base model. The reduction is more significant for lower-rank matrices. For instance, in the LLaMA-3-70B dubey2024llama , all “Key” projection layers see a reduction of 49%. The results in Table 4 validate that QLoRA, discussed in Section 4, does not reduce quantization error. In contrast, PiSSA significantly outperforms LoftQ in reducing quantization error, as further discussed in Appendix H.
@@
The difference between QPiSSA and PiSSA is the quantization of the residual model to 4 bits. As introduced in Section 4, the residual model has less influence on the optimal direction compared with the principal adapter, which is the same in both QPiSSA and PiSSA. Therefore, besides reducing the quantization error, we expect QPiSSA to also converge faster than QLoRA and LoftQ. We train LLaMA 3-8B using LoRA/QLoRA, PiSSA/QPiSSA, LoftQ, and full fine-tuning on MetaMathQA-395K for 3 epochs, recording the loss, gradient norm, and accuracy on GSM8K, as shown in Figure 5.
@@ 5.4 Experiments Across Various Sizes and Types of Models
In this section, we compare (Q)PiSSA and (Q)LoRA across 9 models, ranging from 7-70B parameters, including LLaMA 2-7/13B touvron2023llama , LLaMA-3-8/70B dubey2024llama , Mistral-7B jiang2023mistral , Gemma-7B team2024gemma , and Qwen1.5-7B qwen , Yi-1.5-34B young2024yi and MoE models: DeepSeek-MoE-16B dai2024deepseekmoe and Mixtral-8x7B jiang2024mixtral . These models were fine-tuned on the MetaMathQA-100K and CodeFeedback-100K dataset and evaluated on the GSM8K and HumanEval. DeepSeek-MoE-16B, Mixtral-8x7B, Yi-1.5-34B, and LLaMA-3-70B were fine-tuned with QPiSSA and QLoRA, while the other models were using PiSSA and LoRA. From Figure 6, (Q)PiSSA, compared to (Q)LoRA, shows improved accuracy across various sizes and types of models, demonstrating its consistent advantage over (Q)LoRA.
@@ 5.5 Experiments on Various Ranks
This section explores the impact of incrementally increasing the rank of PiSSA/QPiSSA and LoRA/QLoRA from 1 to 128, aiming to determine whether PiSSA/QPiSSA consistently outperforms LoRA/QLoRA under different ranks. The training is conducted using the MetaMathQA-100K dataset for 1 epoch, while the validation is performed on the GSM8K and MATH datasets. The outcomes of these experiments are depicted in Figure 7, with additional results presented in Appendix K.
@@ 6 Conclusion
This paper presents a PEFT technique that applies singular value decomposition (SVD) to the weight matrix of pre-trained models. The principal components obtained from the SVD are used to initialize a low-rank adapter named PiSSA, while the residual components are kept frozen, to achieve effective fine-tuning and parameter efficiency simultaneously. Through extensive experiments, we found that PiSSA and its 4-bit quantization version QPiSSA significantly outperform LoRA and QLoRA in both NLG and NLU tasks, across different training steps, various model sizes and types, and under various amount of trainable parameters. PiSSA provides a novel direction for research in PEFT by identifying and fine-tuning the principal components within the model, analogous to slicing and re-baking the richest slice of a pizza. As PiSSA shares the same architecture as LoRA, it can be seamlessly used in existing LoRA pipelines as an efficient alternative initialization method.
@@ 7 Limitation
There are still some questions with PiSSA not addressed in this paper: 1) Besides language models, can PiSSA also be adapted to convolutional layers and enhance the performance of vision tasks? 2) Can PiSSA also benefit from some improvements to LoRA, such as AdaLoRA zhang2023adaptive and DyLoRA valipour2022dylora which adaptively adjust the rank? 3) Can we provide more theoretical explanations for the advantages of PiSSA over LoRA? We are actively exploring these questions. Nevertheless, we are excited to see the huge potential of PiSSA already demonstrated in existing experiments and look forward to more tests and suggestions from the community.
8 Acknowledgements:
This work is supported by the National Key R&D Program of China (2022ZD0160300).
Appendix A Enhancing PiSSA with LoRA Improvement Methods
AdaLoRA introduces three improvements over LoRA:
  • Trainable parameters in AdaLoRA are changed to $A,B$, and $E$. $A$ and $B$ are Gaussian-initialized, and $E$ is a zero-initialized $r$-dimensional vector, making $Adiag(E)B=\Delta W$, similar to singular value decomposition.
  • A regularization loss $|AA^{T}-I|+|B^{T}B-I|$ is used to make $A$ and $B$ gradually orthogonal during training, resembling the SVD of $\Delta W$.
  • An initial large rank is set, and less important E values are gradually masked during training, resulting in different final ranks for each layer, achieving better performance with the same number of parameters.
Despite the extensive use of SVD terms, AdaLoRA does not perform actual SVD on any matrix. In the PEFT domain, terms like low-rank decomposition, and singular value decomposition often appear. They generally refer to products of low-dimensional matrices approximating an ideal $\Delta W$ without actual matrix decomposition. To our knowledge, PiSSA is the first to perform SVD on the original model, fine-tuning the principal component while keeping the residual model frozen.
PiSSA and AdaLoRA represent different improvements to LoRA, making them combinable. Therefore, we additionally improved PiSSA based on AdaLoRA’s three innovations:
  • After extracting the principal singular values and vectors of $W$, we use $S$ as an independent trainable vector instead of multiplying it into $U$ and $V$.
  • Since PiSSA’s $U$ and $V$ are orthogonal at the beginning, maintaining their orthogonality through orthogonal regularization is very easy.
  • Although AdaLoRA claims to dynamically reduce the number of trainable parameters, the initially large number of parameters is not truly pruned, resulting in more parameters being updated during actual training. Therefore, we did not use this improvement.
DoRA adds a learnable magnitude module to LoRA, normalizing $W+AB$ at each update step and multiplying its by the magnitude module. This allows $A,B$ to learn the direction and the magnitude module to learn the magnitude of $\Delta W$. While this approach can improve fine-tuning performance, normalizing $W+AB$ at each step results in slower fine-tuning speeds. In contrast, PiSSA only changes LoRA’s initialization method, matching LoRA in training speed and converging faster, thereby reducing training costs.
Table 5: GSM8K accuracy for LoRA and PiSSA when combined with LoRA improvement methods.
Model Method LoRA+ PiSSA+
Vanilla 71.01±0.199 76.75±0.036
LLaMA-3-8B DoRA 72.38±0.189 77.51±0.257
AdaLoRA 72.31±0.202 78.59±0.199
PiSSA, with its intrinsic principal singular values and orthogonal singular vectors, is very suitable for combination with AdaLoRA. According to Table 5. The performance of the improved PiSSA surpasses all the other methods including PiSSA. From lines 1, and 2 of the table, it is evident that the performance of PiSSA combined with DoRA significantly surpasses that of DoRA alone and also exceeds the performance of PiSSA alone. Taking into account the combination experiments of PiSSA with AdaLoRA, it can be inferred that PiSSA benefits from the enhancement techniques of LoRA, demonstrating the potential of PiSSA when integrated with other methods.
@@ Appendix D Comparison of Fine-Tuning in BF16 and FP32 Precision
Appendix E Reducing Quantization Error through Multiple Iteration of SVD
@@
Appendix F Conductive Experiments on Various SVD Components
Appendix G The Residual Matrices having a Narrower Distribution
To intuitively compare the distribution differences between quantized original and residual models, in Figure 3, we took LLaMA 2-7B’s first Query layer as an example to illustrate the distribution of $W$ and $W_{res}$. However, using only one layer of one model is not statistically significant. In this section, we applied PiSSA initialization to LLaMA-2-7B, Mistral-7B, Gemma-7B, and LLaMA-3-8B, and fit the values in every linear layer with Gaussian distribution and calculated their mu and sigma.
(a) LLaMA-2-7B
(b) Mistral-7B-v0.1
(c) Gemma-7B
(d) LLaMA-3-8B
Figure 9: Comparison of Loss and Ratio to the target A and target B for LoRA and PiSSA across the initial 5 steps.
The results in Figure 9 show that the residual models’ means are closer to 0, and the standard deviations are smaller after PiSSA initialization. Thus, $W^{res}$ indeed has a narrower distribution than W in a statistical sense. Nevertheless, the difference is not as large as that in the first layer after averaging all layers, which we suspect is because middle layers in a model tend to have more even eigenvalue distributions due to redundancy and insufficient training.
Appendix H Comparing the Quantization Error of QLoRA, LoftQ and QPiSSA
@@
Appendix I Combining QPiSSA with Various Quantization Methods
In addition to NF4 quantization, QPiSSA can also be combined with GPTQ and INT8 quantization. We posit that PiSSA effectively reduces quantization error for several reasons:
  • It reduces outlier values;
  • It makes the value distribution more Gaussian-like;
  • It preserves larger values in full precision, thereby narrowing the weight distribution in the quantized portion.
While INT8 also targets the reduction of outlier values (point 1), PiSSA has the potential to enhance this effect. The second point aligns well with NF4, and the third point is crucial as PiSSA uses an adaptor to retain a significant portion of weights in full precision, maintaining the integrity of critical values.
Table 9: Quantization Error and Accuracy for PiSSA Combined with Various Quantization Methods. GPTQ quantizes each row $w$ independently, adjusting one weight at a time while updating all remaining, non-quantized weights. Therefore, the nuclear norm method used for calculating quantization error in the main paper is not applicable to GPTQ. Instead, we measure Perplexity on WikiText-2, where a lower Perplexity indicates reduced quantization error.
Model Dataset Quantization Error GSM8K Accuracy
QLoRA PiSSA QLoRA PiSSA
NF4 324.8 (nuclear norm) 265.8 (nuclear norm) 70.79±0.42 73.76±0.20
LLaMA-3-8B INT8 34.47 (nuclear norm) 28.21 (nuclear norm) 71.68±0.14 76.54±0.32
GPTQ 20.79 (perplexity) 6.23 (perplexity) 70.18±0.42 74.58±0.22
As shown in Table 9, QPiSSA combined with INT8 reduces quantization error by 18.16% on LLaMA-3-8B, and significantly outperforms QLoRA using INT8. Furthermore, in row 3 of the table, the perplexity of LLaMA-3-8B increases to 20.79 after quantization with GPTQ-4bit on the C4 dataset. However, when PiSSA is applied, the perplexity is reduced to 6.23. These results confirm the effectiveness of PiSSA in reducing quantization error, as discussed in the main paper.
Overall, QPiSSA demonstrates a clear advantage over QLoRA when combined with various quantization methods, retaining the fast convergence and superior performance characteristics of PiSSA while minimizing quantization error.
Appendix J Evaluating PiSSA on Mixtral and Gemma with More Training Steps
This is the supplement for Section 5.2. We applied PiSSA, LoRA, and full parameter fine-tuning on the full MetaMathQA-395K dataset using Mistral-7B and Gemma-7B models, training for 3 epochs. Figures 12 and 13 display the training loss, gradient norm, and evaluation accuracy on GSM8K.